40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 249 of 339  Not logged in ELOG logo
ID Date Author Type Category Subject
  4565   Mon Apr 25 12:55:19 2011 JenneBureaucracyComputerswiki?

Quote:

Quote:

40m wiki seems to have been down for quite a while now but I can't see any info in the elog about it.  Is there some ongoing problem?

 There was an email from Dave Barker about this.  They had to reorganize the DNS at LHO.  The URL that should be used is: http://blue.ligo-wa.caltech.edu:8000/40m

 Nope.  I know I had the right address, and it was down for me too all weekend.  It's better now though.  blue.ligo-wa.caltech.edu was up though.

  4564   Mon Apr 25 11:58:37 2011 JamieBureaucracyComputerswiki?

Quote:

40m wiki seems to have been down for quite a while now but I can't see any info in the elog about it.  Is there some ongoing problem?

 There was an email from Dave Barker about this.  They had to reorganize the DNS at LHO.  The URL that should be used is: http://blue.ligo-wa.caltech.edu:8000/40m

  4563   Mon Apr 25 11:23:41 2011 AlastairBureaucracyComputerswiki?

40m wiki seems to have been down for quite a while now but I can't see any info in the elog about it.  Is there some ongoing problem?

119427.strip.gif

  4562   Sun Apr 24 21:37:40 2011 kiwamuUpdateIOOreview of triple resonant EOM : model looks fine

To design a new resonant EOM box I started reviewing the prototype that I've built.

As a part of reviewing I checked an important thing that I haven't carefully done so far :

I compared the measured input impedance with that of predicted from a circuit model. I found that they show a good agreement.

So I am now confident that we can predict / design a new circuit performance.


* * * (input impedance) * * *

 Performance of a resonant circuit is close related to its input impedance and hence, in other words, determined by the input impedance.

Therefore an investigation of input impedance is a way to check the performance of a circuit. That's why I always use impedance for checking the circuit.

The plot below is a comparison of input impedance for the measured one and one predicted from a model. They show a good agreement.

(Note that the input impedance is supposed to have 50 Ohm peaks at 11, 29.5 and 55 MHz.)
Input_impedance.png

 
* * * (circuit model) * * *

To make the things simpler I assume the following three conditions in my model:

 1. inductor's loss is dominated by its DC resistance (DCR)

 2. capacitor's loss is characterized only by Q-value

 3. Transformer's loss is dominated by DCR and its leakage inductance

All the parameters are quoted from either datasheet or my measurement. The model I am using is depicted in the schematic below.

Basically the Q-vaules for the capacitors that I used are quite low. I think higher Q capacitors will improve the performance and bring them to more 50 Ohm.

EOMcircuitmodel.png
 

  4561   Fri Apr 22 12:07:38 2011 josephb, steveUpdateCDSRemoved hanging D-sub to SCSI in 1X2

Problem:

Way back, Jay had D-sub to SCSI adapters made to adapt our existing Sander box AA filters to the new SCSI based IO chassis.  However, these did not fit inside the box.

At the time, we simply left the cards outside hanging, which was a hack and needed to be replaced.

Solution:

Steve modified a black AA filter box so that it could fit the D-sub to SCSI adapter board on it, plus strain relief the SCSI cable, rather than let it hang.  The back of the box was cut, and an extending piece of metal attached to the bottom of the box.  The adapter board was screwed into the box, the SCSI plugged in, then the SCSI cable is clamped to the extending metal as well.

This modification will be propagated to the 3 remaining AA filter boards using the D-sub to SCSI adapter.

  4560   Fri Apr 22 11:08:50 2011 kiwamuUpdateLSCdemod board AS11 : amplitude imbalance

Amplitude imbalance between I and Q in a demod board, AS11, with the new 90 deg splitter was measured.

It shows roughly 10% amplitude imbalance when the LO power is in a range from 0 to 5 dBm. Not so bad.

 

  With the handmade coil there used to be a huge imbalance (either I or Q goes to zero volt while the other keeps about 1 V rms) as the LO power decreases.

But with the new 90 deg splitter now there are no more such a huge imbalance.

The remaining 10 % imbalance possibly comes from the fact that we are using ERA-5 in each I and Q path. They may have such gain imbalance of 10%.

We should check the ERA-5 gains so that we can confidently say ERA-5 causes the amplitude imbalance.

Then our plan replacing the ERA-5s (see here) will sound more reasonable.

IQamplitude.png

 

Quote from #4555

The new 90 deg splitter works better.

 I will also measure amplitude unbalances between I and Q.

 

  4559   Fri Apr 22 10:28:22 2011 ranaUpdateRF SystemRF Source Harmonics
You should be able to resolve the other harmonics by decreasing the IF BW or RBW on the analyzer. Even though
they're OK, its useful to have the final measurement of all of them in some kinds of physical units (like dBm, but
not dBm/Hz or dB or dBcubits).
  4558   Fri Apr 22 09:25:43 2011 SureshUpdateRF SystemRF Source: Temperature sensor relocated

RF Amp operating temperature

Earlier measurement reported by Alberto in LIGO-T10004-61-v1 based on the LM34 temperature sensor were lower than that shown by placing a calibrated thermocouple sensor directly on the heat sink by about 5deg C. The difference probably arose because the LM34 was located on a separate free-hanging copper sheet attached to the RF Amp by a single screw, resulting in a gradient across the copper strip.   I tried to move the LM34 which was glued down, but broke the leads in the process.  I then replaced it with another one mounted much closer to the heat sink and held it down with a copper-strip clamp.  There is no glue involved and there is heatsink compound between the flat surface of the LM34 and the heatsink.  Picture attached. 

  The picture also shows the new filters which have been put in place to reduce the harmonics.  Note that the SBP-10.7 which was to go on the 11 MHz Demod output is located much farther upsteam due to space constraints.

P4220056.JPG

  4557   Fri Apr 22 09:05:53 2011 SureshUpdateRF SystemRF Source Harmonics
As seen in the previous measurement the first harmonic of both the 11 MHz and 55 MHz outputs are about 30dB
higher than desired.  In an attempt to attenuate these and higher harmonics I introduced SBP-10.7 filters into
the 11MHz outputs and SLP-50 filters into the 55 MHz outputs.
Then I measured the height of the harmonics again and found that they were suppressed as expected.  Now harmonic
at 22 MHz is 58dB lower than the 11 MHz fundamental.  And the 110 MHz is lower by 55 dB compared to the 55 MHz
fundamental.  None of the higher harmonics are seen => they are below 70dB

SLP-50 has an insertion loss(IL) of 4.65 dB and Return Loss(RL) of 3dB.  It would be better to use SBP-60
(IL=1.4 dB and RL=23dB)

The filter on the 11 MHz lines is okay. The SBP-10.7 has IL=0.6 dB and RL=23 dB.
  4556   Fri Apr 22 02:10:53 2011 ZachUpdateelogrestarted

Restarted the elog with the script.

  4555   Thu Apr 21 21:46:22 2011 kiwamuUpdateLSCdemod board : new 90 deg splitter

A less LO power dependence on the relative phase was found. The new 90 deg splitter works better.

From -3 dBm to 10 dBm in LO power, the relative phase is within 90 +/- 5 deg.

As a comparison I plot the phase that I measured when the handmade coil had been there (green curve in the plot).

relativephase.png

 

 I will also measure amplitude unbalances between I and Q.

Quote from #4554

A 90 degree splitter, PSCQ-2-51W, has arrived today and I installed it on a demod board called AS11.

Results of the I-Q phase measurement with the new splitter will be reported soon.

 

 

  4554   Thu Apr 21 21:24:41 2011 kiwamuUpdateLSCdemod board : new 90 deg splitter

A new 90 degree splitter, PSCQ-2-51W, has arrived today and I installed it on a demod board called AS11.

Results of the I-Q phase measurement with the new splitter will be reported soon.

 

 * Picture 1 = before removal of the handmade coil

 * Picture 2 = after removal of the coil and the associated capacitors

 * Picture 3 = after soldering PSCQ-2-51-W

DSC_2949_ss.jpg

DSC_2951.JPG

DSC_2952_ss.jpg

Quote from #4358

 First of all we will replace the home-made 90 degree splitter (see this entry) by a commercial splitter, PSCQ-2-51-W+ from Mini circuit. This is the step 1 basically.

  4553   Thu Apr 21 15:20:46 2011 steveUpdatePEMPSL enclosure work today

Quote:

The east side window guides  will be replaced by one long U-channel. There will be drilling into 2x2 steel frame. It should be done by 2pm today

This should remove the jerking motion of windows hitting the individual guides.

 The installation went smoothly. There will be no more banging the doors into  guides on the east side.

Atm2 showing the west side as is today

Attachment 1: P1070558.JPG
P1070558.JPG
Attachment 2: P1070560.JPG
P1070560.JPG
  4552   Thu Apr 21 15:03:29 2011 steveUpdateComputersjuction board finds home

The anti aliasing box was opened up at the back to accommodate the junction board and the SCSI cable towards the ADC. Aluminum plate was attached to the bottom to hold the strain relief clamp.

Three more hanging junction cards will be replaced in this manner.

Attachment 1: P1070570.JPG
P1070570.JPG
Attachment 2: P1070568.JPG
P1070568.JPG
  4551   Thu Apr 21 14:39:43 2011 steveUpdateRF Systemnew strain relieved N connectors at AP

New right angle PVC, 2 x 2 x  1/4" installed at the AP table to strain relief the 1/4" spiral corrugated RF coaxes.

Attachment 1: P1070562.JPG
P1070562.JPG
Attachment 2: P1070564.JPG
P1070564.JPG
  4550   Thu Apr 21 08:26:42 2011 steveUpdatePEMPSL enclosure work today

The east side window guides  will be replaced by one long U-channel. There will be drilling into 2x2 steel frame. It should be done by 2pm today

This should remove the jerking motion of windows hitting the individual guides.

  4549   Wed Apr 20 23:20:49 2011 jamieSummaryComputersinstallation of CDS tools on pianosa

This is an overview of how I got (almost) all the CDS tools running on pianosa, the new Ubuntu 10.04 control room work station.

This is machine is experiment in minimizing the amount of custom configuration and source code compiling. I am attempting to install as many tools as possible from existing packages in

available packages

I was able to install a number of packages directly from the ubuntu archives, including fftw, grace, and ROOT:

apt-get install \
libfftw3-dev \
grace \
root-system

LSCSOFT

I installed all needed LSCSOFT packages (framecpp, libframe, metaio) from the well-maintained UWM LSCSOFT repository.

$ cat /etc/apt/sources.list.d/lscsoft.list
deb http://www.lsc-group.phys.uwm.edu/daswg/download/software/debian/ squeeze
deb-src http://www.lsc-group.phys.uwm.edu/daswg/download/software/debian/ squeeze contrib
sudo apt-get install lscsoft-archive-keyring
sudo apt-get update
sudo apt-get install ldas-tools-framecpp-dev libframe-dev libmetaio-dev lscsoft-user-en

You then need to source /opt/lscsoft/lscsoft-user-env.sh to use these packages.

EPICS

There actually appear to be a couple of projects that are trying to provide debs of EPICS. I was able to actually get epics working from one of them, but it didn't include some of the other needed packages (such as MEDM and BURT) so I fell back to using Keith's pre-build binary tarball.

Prereqs:

apt-get install \
libmotif-dev \
libxt-dev \
libxmu-dev \
libxprintutil-dev \
libxpm-dev \
libz-dev \
libxaw7-dev \
libpng-dev \
libgd2-xpm-dev \
libbz2-dev \
libssl-dev \
liblapack-dev \
gfortran

Pulled Keith's prebuild binary:

cd /ligo/apps
wget https://llocds.ligo-la.caltech.edu/daq/software/binary/apps/ubuntu/epics-3.14.10-ubuntu.tar.gz
tar zxf epics-3.14.10-ubuntu.tar.gz

GDS

I built GDS from svn, after I fixed some broken stuff [0]:

cd ~controls/src/gds
svn co https://redoubt.ligo-wa.caltech.edu/svn/gds/trunk
cd trunk
#fixed broken stuff [0]
source /opt/lscsoft/lscsoft-user-env.sh
./bootstrap
export GDSBUILD=online
export ROOTSYS=/usr
./configure --prefix=/ligo/apps/gds --enable-only-dtt --with-epics=/ligo/apps/epics-3.14.10
make
make install

dataviewer

I installed dataviewer from source:

cd ~controls/src/advLigoRTS
svn co https://redoubt.ligo-wa.caltech.edu/svn/advLigoRTS/trunk
cd trunk/src/dv
#fix stupid makefile /opt/rtapps --> /ligo/apps
make
make install

I found that the actual dataviewer wrapper script was also broken, so I made a new one:

$ cat /ligo/apps/dv/dataviewer
#!/bin/bash
export DVPATH=/ligo/apps/dv
ID=$$
DCDIR=/tmp/${ID}DC
mkdir $DCDIR
trap "rm -rf $DCDIR" EXIT
$DVPATH/dc3 -s ${NDSSERVER} -a $ID -b $DVPATH "$@"

environment

Finally, I made a environment definer file:

$ cat /ligo/apps/cds-user-env.sh
# source the lscsoft environment
. /opt/lscsoft/lscsoft-user-env.sh

# source the gds environment
. /ligo/apps/gds/etc/gds-user-env.sh

# special local epics setup
EPICS=/ligo/apps/epics
export LD_LIBRARY_PATH=${EPICS}/base/lib/linux-x86_64:$LD_LIBRARY_PATH
export LD_LIBRARY_PATH=${EPICS}/extensions/lib/linux-x86_64:$LD_LIBRARY_PATH
export LD_LIBRARY_PATH=${EPICS}/modules/seq/lib/linux-x86_64:$LD_LIBRARY_PATH
export PATH=${EPICS}/base/bin/linux-x86_64:$PATH
export PATH=${EPICS}/extensions/bin/linux-x86_64:$PATH
export PATH=${EPICS}/modules/seq/bin/linux-x86_64:$PATH

# dataviewer path
export PATH=/ligo/apps/dv:${PATH}

# specify the NDS server
export NDSSERVER=fb

[0] GDS was not compiling, because of what looked like bugs. I'm not sure why I'm the first person to catch these things. Stricter compiler?

To fix the following compile error:

TLGExport.cc:1337: error: ‘atoi’ was not declared in this scope

I made the following patch:

Index: /home/controls/src/gds/trunk/GUI/dttview/TLGExport.cc
===================================================================
--- /home/controls/src/gds/trunk/GUI/dttview/TLGExport.cc (revision 6423)
+++ /home/controls/src/gds/trunk/GUI/dttview/TLGExport.cc (working copy)
@@ -31,6 +31,7 @@
#include <iomanip>

#include <string.h>

#include <strings.h>

+#include <stdlib.h>


namespace ligogui {
using namespace std;

To fix the following compile error:

TLGPrint.cc:264: error: call of overloaded ‘abs(Int_t&)’ is ambiguous

I made the following patch:

Index: /home/controls/src/gds/trunk/GUI/dttview/TLGPrint.cc
===================================================================
--- /home/controls/src/gds/trunk/GUI/dttview/TLGPrint.cc (revision 6423)
+++ /home/controls/src/gds/trunk/GUI/dttview/TLGPrint.cc (working copy)
@@ -22,6 +22,7 @@
#include <fstream>

#include <map>
#include <cmath>

+#include <cstdlib>


namespace ligogui {
using namespace std;

  4548   Wed Apr 20 22:29:07 2011 sureshUpdateRF SystemPlan for LSC rack

The suggested layout of the 1Y2 Rack is shown below.

To simplify the wiring, I have largely kept demod boards with the same same LO frequency close to each other. 

The Heliax cables land on the top and bottom of the of subracks.  These are currently flexible plastic sheets.  Steve has agreed to replace them with something more rigid.  It would be good to have eight N-type connectors on the top and eight  at the bottom.  As  demod boards occur in sets of eight per subrack.  So it would be convenient if the 11 and 55 Mhz Heliax cables land on the top and the rest at the bottom.  In the layout I have shown the current situation. 

The LO signals to the boards come from the RF Distribution box and this is kept in the middle so that cables to both the subracks can be kept short.

The outputs of the AA filter boards from both subracks  have to be connected to the SCSI Interface board with a twisted pair ribbon cable. 

1Y2_Rack_Layout.png

  4547   Wed Apr 20 21:53:01 2011 SureshConfigurationRF SystemRF system: Stray heliax cable

We found a stray unused heliax cable running from the LSC rack 1Y2 to a point between the cabinets 1X3 and 1X4. This cable will need to be redirected to the AS table in the new scheme.   It is labled C1LSC-PD5  The current situation has been updated as seen in the layout below

rogue_cable_1.png

Attachment 1: rogue_cable_1.png
rogue_cable_1.png
  4546   Wed Apr 20 20:42:55 2011 kiwamuUpdateVIDEOtoday's video session

[Steve / Suresh / Kiwamu]

90 % of unused video cables have been removed.

Still a couple of video cables are floating around the video MUX. They will be removed in the next week's session.

DSC_2931_ss.jpg

DSC_2938_ss.jpg

 

  4545   Wed Apr 20 11:02:18 2011 josephbUpdateCDSMEDM screens and Front Ends updated to new Matrices
We simply didn't any matrices larger than 16x16. If we had, than that matrix would not have worked properly since the beginning.

Quote:

Just a curiosity:

I just wonder how you have distingushed the difference between _111 and _111.

They are equivalent alone themselves. Have you looked at the contexts of the lines?
Or you just did not have the larger matrix than 16x16, did you?

 

  4544   Tue Apr 19 17:34:02 2011 KojiUpdateCDSMEDM screens and Front Ends updated to new Matrices

Just a curiosity:

I just wonder how you have distingushed the difference between _111 and _111.

They are equivalent alone themselves. Have you looked at the contexts of the lines?
Or you just did not have the larger matrix than 16x16, did you?

  4543   Tue Apr 19 15:48:43 2011 josephbUpdateCDSMEDM screens and Front Ends updated to new Matrices

Problem:

The original matrix naming conventions for the front ends was broken.  It used _11, _12,...,_1e, _1f, _110, _111 and so forth.  The code was changes to use _1_1, _1_2,...,_1_16,_1_17, and so on.

In addition the matrix of filter banks was modified to use the same naming convetion (instead of starting at zero, it now start with one).

Work Done:

I rebuilt all the models, and restarted them all.

I wrote a simple script to modify the burt restore files to have the correct names for all the stored matrix values.

I also modified all the suspension screens, by modifying the default screens in /opt/rtcds/caltech/c1/medm/master/

The C1SUS, C1SCX, C1SPX, C1SCY, C1SPY, and C1MCS models had their foton filter files modified to put filters into the newly changed named filters

  4542   Mon Apr 18 21:14:53 2011 JamieConfigurationComputersnew control room machine: pianosa
Also, op440m's Sun monitor did not work well with pianosa, so I'm lending pianosa my HP monitor until we can get a suitable replacement.
  4541   Mon Apr 18 21:09:45 2011 JamieConfigurationComputersnew control room machine: pianosa

I've just installed the new control room machine: "pianosa".   It is a replacement for the old sun machine "op440m" [0].

Hardware:

  • dual dual-core Intel Core i7-2600 CPU @ 3.40GH, hyperthreaded to provide 8 effective cores
  • 16G memory (4x 4G dimms)
  • nVidia GF108 GeForce GT 430

It's now running Ubuntu 10.04 LTS 64bit.  Unfortunately, the default 10.04 kernel is 2.6.32, which does not support pianosa's apparently very new network adapter, which is (from lspci):

00:19.0 Ethernet controller: Intel Corporation 82579LM Gigabit Network Connection (rev 04)

To get around this I temporarily added a PCI nic so that I could get on the network.  I then added the Ubuntu kernel team PPA archive and installed linux-image-2.6.38-2, which is new enough to have the needed network driver, but not completely bleeding edge:

sudo add-apt-repository ppa:kernel-ppa/ppa
sudo apt-get update
sudo apt-get install linux-image-2.6.38-2-generic linux-headers-2.6.38-2-generic

Once the built in nic was working I removed the temporary one.  Everything seems to be working fine now.

I have not yet done any configuration to integrate pianosa into the CDS network.  I'll do that tomorrow.

[0] op44m has been moved into the control room rack next to linux1, in headless mode.  If there is still a need to run scripts that only run on solaris, op440m can still be accessed via ssh as normal.  Hopefully we can fully decommission this machine soon.

[1] https://launchpad.net/~kernel-ppa/+archive/ppa

  4540   Mon Apr 18 17:47:41 2011 kiwamuConfigurationLSCLSC rack's ADC cabling

To understand the situation of the ADC cabling at the LSC rack I looked around the rack and the cables.

The final goal of this investigation is to have nice and noise less cables for the ADCs (i.e. non-ribbon cable)

Here is just a report about the current cabling.

 

(current configuration)

At the moment there is only one ribbon-twisted cable going from 1Y2 to 1Y3. (We are supposed to have 4 cables).

At the 1Y2 rack the cable is connected to an AA board with a 40 pin female IDC connector.

At the 1Y3 rack the cable is connected to an ADC board with a 37 pin female D-sub connector.

The ribbon cable is 28AWG with 0.05" conductor spacing and has 25 twisted pairs (50 wires).

LSCrack.png

 

(things to be done)

 - searching for a twisted-shielded cable which can nicely fits to the 40 pin IDC and 37 pin D-sub connectors.

 - estimating how long cable we need and getting the quote from a vendor.

 - designing a strain relief support

  4539   Mon Apr 18 14:11:44 2011 kiwamuUpdateLSCRF status

 We will make them all green !!

 RF_Work_Status.png

Again, all the files are available in the svn.

https://nodus.ligo.caltech.edu:30889/svn/trunk/suresh/40m_RF_upgrade/

  4538   Mon Apr 18 13:05:57 2011 kiwamuSummaryLSCdemod board modification

Here is the idea how we upgrade the demodulation boards.

Basically we go ahead with two steps as depicted in the cartoon diagram below.

Once we finish the first step of upgrade, the board will be ready to install although the circuit won't be awesome in terms of noise performance.

 

demod_board.png

 

* * * (details) * * *

 First of all we will replace the home-made 90 degree splitter (see this entry) by a commercial splitter, PSCQ-2-51-W+ from Mini circuit. This is the step 1 basically.

At this point the boards will be ready to use in principle. I asked Steve to get three 90 degree splitters so that we can have at least three demodulators for the dual-recycled Michelson locking.

If they work very fine we will buy some more 90 degree splitters for full locking.

While we try to lock the dual-recycled Michelson once we will get a Cougar amplifier, remove all ERA-5s and install it such that we don't have to gain up and down in the circuit. This is the last step.

  4537   Sat Apr 16 02:00:14 2011 ranaUpdateLSC11MHz demod board : 90 degree splitter

One way to avoid some of the bad stuff in there is to take the 1 dBm input and amplify it to ~21 dBm before splitting and sending in to the Level 17 mixers.

One way to do this is by using the A3CP6025 from Teledyne-Cougar. Its an SMA connectorized amp which can put out 25 dBm and has a gain of 24 dB. We can just glue it onto the demod boards. Then we can remove the ERA-5 amplifiers and just use the broadband splitter as Kiwamu mentioned.

  4536   Fri Apr 15 22:57:38 2011 Aidan, JamieConfigurationLSCAS PD and Camera installed

AS port ITMX YAW  range where AS beam was visible = [-1.505, -1.225] - these extrema put the beam just outside of some aperture in the system -> set ITMX YAW to -1.365

ITMX PITCH range = [-0.7707, -0.9707] -> set to ITMX PITCH to -0.8707

Quote:

I spent some time tracking down the AS beam which had vanished from the AP table. Eventually, by dramatically mis-aligning SRM, PRM and ITMY, returning BS to its Jan 1st PITCH and YAW values and tweaking the ITMX alignment [actual values to follow], I was able to get an AS beam out onto the AP table. I verified that it was the prompt reflection off ITMX by watching it move as I changed the YAW of that optic and watching it stay stationary as I changed the YAW of ITMY.

Jamie and I then steered the beam through a 2" PLCX-50.8-360.6 lens and placed the RF PD (AS55) at the focus. Additionally, we installed the AS camera to observe the leakage field through a Y1S steering mirror (as shown in the attached diagram).

Currently the PD has power but the RF and DC outputs are not connected to anything at the moment.

Atm 2 by Steve

 

 

 

  4534   Fri Apr 15 22:54:20 2011 Aidan, BryanUpdateGreen LockingBeat note amplitude on Vertex PD

I was investigating the beat note amplitude on the vertex PD again yesterday. The incident power on the PD was 150uW in the PSL green beam and 700uW in the X-ARM green beam. With perfect overlap and a transimpedance of 240, I expected to get a beat note signal of around 25mV or -19dBm. Instead, the size was -57dBm. Bryan and I adjusted the alignment of the green PSL beam to try and improve the mode overlap but we couldn't do much better than about -50dBm. (The noise floor of the PD is around -65dBm).

When we projected the beams to the wall of the enclosure, the xarm beam was 2 to 3x as large as the PSL green beam, indicating that the beam size and/or curvatures on the PD were less than ideal. There is a telescope that the XARM beam goes through just before it gets to the PD. I mounted the second lens in this telescope on a longitudinal translation stage. With some finagling of the position of that lens we were able to improve the beatnote signal strength to -41dBm.

Obviously the ideal solution would be to measure the beam size and RoC of the PSL beam and XARM beams and then design a telescope that would match them as precisely as possible because there's still another 20dB signal strength to be gained.

 

  4533   Fri Apr 15 15:15:08 2011 kiwamuUpdateLSC11MHz demod board : 90 degree splitter

[Rana, Koji, Kiwamu]

 Moreover the amplitude of the I and Q signals are highly unbalanced, depending on the LO power again.

This implies the coil for a 90 degree splitting won't work at 11 MHz since the coil is home made and used to be designed for a specific frequency (i.g. 24.5 MHz).

We decided to use a Mini circuit 90 deg splitter instead. Steve will order few of them soon and we will test it out.

Quote:

During checking the 11MHz demod boards I found that the I-Q relative phase showed funny LO power dependence.

It is now under investigation.

 

  4532   Fri Apr 15 13:43:23 2011 BryanConfigurationGreen LockingI leave you with these messages...

Y-end PDH electronics.

The transfer function of the Y-end universal PDH box:

Y_End_Electronics_TF.png

 

  4531   Fri Apr 15 13:40:00 2011 Larisa ThorneUpdateElectronicsLow pass filter for X arm laser temperature control, second try

Plotting the data points yielded by the spec analyzer of my first LPF yielded a result that was not expected: the desired cutoff frequency wasn't achieved because of some extra 100k resistance that wasn't taken into consideration. (see  here ). I have redrawn the Bode graphs for this configuration so that it is easier to see that it is wrong (first attachment)

 

After some calculation adjustments, it was found that the capacitor value could remain at 10uF, but the resistance needed to be changed to 100k to maintain a gain of 0.5 and critical frequency at 0.1Hz. Second attachment is the Bode graph that results from this configuration.

 

Note: Bode graphs are both in Log-Linear scales (Wikipedia said so)

 

Attachment 1: Bode2.jpeg
Bode2.jpeg
Attachment 2: Bode100k.jpg
Bode100k.jpg
  4530   Fri Apr 15 12:17:39 2011 kiwamuUpdateLSC11MHz demod board : funny I-Q phase

During checking the 11MHz demod boards I found that the I-Q relative phase showed funny LO power dependence.

It is now under investigation.

relativephase.png

 In the plot above the green curve represents the I-Q phase of a 11MHz demod board (see here).

It showed a strong dependence on the LO power and it changes from -60 deg to -130 deg as the LO power changes.

This is not a good situation because any power modulation on the LO will cause a phase jitter.

For a comparison I also took I-Q relative phase of a 33MHz demod board, which hasn't been modified recently.

 It shows a nice flat curve up to 5 dBm although it looks like my rough measurement adds a systematic error of about -5 deg.

 

 - to do -

* check RF power in every point of LO path on the circuit

* check if there is saturation by looking at wave forms.

  4529   Fri Apr 15 02:30:24 2011 KojiHowToVIDEOVideo switching tip

I have made a small python script to handle the video matrix.

It is too far from the perfection, but I release it as it is already useful in some extent.

The script is in the /cvs/cds/rtcds/caltech/c1/scripts/general directory.

usage:

videoswitch.py in_ch_name out_ch_name

in_ch_name is one of the followings

MC2F, IFOPO, OMCR, FI, AS_Spare, ITMYF, ITMXF, ETMYF, ETMXF,
PMCR, RCR, RCT, PSL_Spare, PMCT, ETMXT, MC2T, POP, IMCR, REFL,
MC1F, SRMF, AS, ETMYT, PRM, OMCT, Quad1, Quad2, Quad3

out_ch_name is one of the followings

Mon1, Mon2, Mon3, Mon4, Mon5, Mon6, Mon7,
ETMY, MC1, PSL1, PSL2, ETMX, MC2, CRT9,CRT10,Projector,
Quad1_1, Quad1_2, Quad1_3, Quad1_4,
Quad2_1, Quad2_2, Quad2_3, Quad2_4,

Quad3_1, Quad3_2, Quad3_3, Quad3_4

  4528   Fri Apr 15 02:18:50 2011 KojiUpdateLSCBunch of RF cables removed

While Kiwamu was working on the RF cabling at the LSC rack, I removed 80% of SMA cables which were not connected anywhere.
The rack is cleaner now, but not perfect yet. We need patch panels/strain relieving for heliaxes, cleaning up of the RF/LO cables, etc.

  4527   Fri Apr 15 02:17:18 2011 kiwamuUpdateLSCMichelson locked

[Koji / Kiwamu]

The Michelson was locked with the new LSC realtime code.

 

 

(what we did)

 --  Fine alignment of the Michelson, including PZTs, BS and ITMY.

  Since the X arm has been nicely aligned we intentionally avoided touching ITMX. The IR beam now is hitting the center of both end mirrors.

  At the end we lost X arm's resonance for IR. This probably means the PZTs need more careful alignments.

 

-- Signal acquisition

 We replaced the RFPD (AS55) that Aidan and Jamie nicely installed by POY11 because we haven't yet  installed a 55MHz RF source.

The maximum DC voltage from the PD went to about 50 mV after aligning steering mirrors on the AP table.

The RF signal from the PD is transferred by a heliax cable which has been labeled 'REFL33'.

Then the RF signal is demodulated at a demodulation board 'AS11', which is one of the demodulation boards that Suresh recently modified.

Although we haven't fully characterized the demod board the I and Q signal looked healthy.

Finally the demod signals go to ADC_0_3 and ADC_0_4 which are the third and fourth channel.

They finally show up in REFL33 path in the digital world.

 

-- Control

 With the new LSC code we fedback the signal to BS. We put anti-whitening filters in the I and Q input filter banks.

We found that dataviewer didn't show correct channels, for example C1LSC_NREFL33I showed just ADC noise and C1LSC_NREFL33Q showed NREFL_33I.

Due to this fact we gave up adjusting the digital phase rotation and decided to use only the I-phase signal.

Applying a 1000:10 filter gave us a moderate lock of the Michelson. The gain was -100 in C1LSC_MICH_GAIN and this gave us the UGF of about 300 Hz.

 Note that during the locking both ETMs were intentionally misaligned in order not to have Fabry-Perot fringes.

  4526   Thu Apr 14 19:05:17 2011 KojiUpdateLSCNew (temporary) LSC screen

[Jenne Koji]

The PD signals are transmitted to the suspension now.

The trigger thresholds were set to -1. This means the triggers are always on.

Attachment 1: temporary_LSC_screen.png
temporary_LSC_screen.png
  4525   Thu Apr 14 17:45:59 2011 BryanConfigurationGreen LockingI leave you with these messages...

OK… the Y-arm may be locked with green light, which was the goal, and this is all good but it's not yet awesome. Awesome would be locked and aligned properly and quiet and optimised. So...  in order to assist in increasing the awesome-osity, here are a few stream-of-consciousness thoughts and stuff I've noticed and haven't had time to fix/investigate or have otherwise had pointed out to me that may help...

 

Firstly, the beam is not aligned down the centre of the cavity. It's pretty good horizontally, but vertically it's too low by about 3/4->1cm on ETMY. The mirrors steering the beam into the cavity have no more vertical range left, so in order to get the beam higher the final two mirrors will have to be adjusted on the bench. Adding another mirror to create a square will give more range AND there will be less light lost due to off 45degree incident angles. When I tried this before I couldn't get the beam to return through the Faraday, but now the cavity is properly aligned this should not be a problem.

 

A side note on alignment - while setting cameras and viewports and things up, Steve noticed that one of the cables to one of the coils (UL) passes behind the ETMY. One of the biggest problems in getting the beam into the system to begin with was missing this cable. It doesn't fall directly into the beam path if the beam is well aligned to the cavity, but for initial alignment it obscures the beam - this may be a problem later for IR alignment.

 

Next, the final lambda/2 waveplate is not yet in the beam. This will only become a problem when it comes to beating the beams together at the vertex, but it WILL be a problem. Remember to put it in before trying to extract signals for full LSC cavity locking.

 

Speaking of components and suchlike things, the equipment for the green work was originally stored in 3 plastic boxes which were stored near the end of the X-arm. These boxes, minus the components now used to set up the Y-end, are now similarly stored near the end of the Y-arm.

 

Mechanical shutter - one needs to be installed on the Y-end just like the X-end. Wasn't necessary for initial locking, but necessary for remote control of the green light on/off.

 

Other control… the Universal PDH box isn't hooked up to the computers. Connections and such should be identical to the X-arm set-up, but someone who knows what they're doing should hook things up appropriately.

 

More control - haven't had a chance to optimise the locking and stability so the locking loop, while it appears to be fairly robust, isn't as quiet as we would like. There appears to be more AM coupling than we initially thought based on the Lightwave AM/PM measurements from before. It took a bit of fiddling with the modulation frequency to find a quiet point where the apparent AM effects don't prevent locking. 279kHz is the best point I've found so far. There is still a DC offset component in the feedback that prevents the gain being turned up - unity gain appears limited to about 1kHz maximum. Not sure whether this is due to an offset in the demod signal or from something in the electronics and haven't had time left to check it out properly yet. Again, be aware this may come back to bite you later.

 

Follow the bouncing spot - the Y-arm suspensions haven't been optimised for damping. I did a little bit of fiddling, but it definitely needs more work. I've roughly aligned the ETMY oplev since that seems to be the mass that's bouncing about most but a bit of work might not go amiss before trusting it to damp anything.

 

Think that's about all that springs to mind for now…

 

Thanks to everyone at the 40m lab for helping at various times and answering daft questions, like "Where do you keep your screwdrivers?" or "If I were a spectrum analyser, where would I be?" - it's been most enjoyable!

 
  4524   Thu Apr 14 12:57:15 2011 josephbUpdateCDSRFM network happy again

[Joe, Alex]

Problem Symptoms:

There were red lights on the status screen indicating RFM errors for the c1scy, c1mcs and c1rfm processes.

The c1iscey, c1sus machines were receiving data sent over the RFM network from the c1ioo computer with a bad time stamp, a few cycles too late.  The c1iscex computer was receiving data from c1ioo fine.

Problem:

The c1iscex RFM card had gotten into a bad state and was somehow slowing things down/corrupting data.  It didn't affect itself, but due to the loop topology was messing everyone else up.  Basically the only one who wasn't throwing an error was the culprit.

Solution:

Hard power cycling the c1iscex computer reset the RFM card and fixed the problem.

  4523   Thu Apr 14 01:03:43 2011 KojiUpdateLSCLSC Campaign ~ Status

Target: To lock the Michelson with the new RF/LSC


Status

RF generation box: READY - already ready to go to the IOO rack. (Suresh)

RF distribution box: In Progress - the internal components are to be connected. (13th evening - Suresh)

Placing PD and CCD: Done - PD and CCD on the AP table (13th Afternoon - Aidan, Larisa with supervision of Kiwamu)

Cabling1: Done - PD signal AP table to the demodulator (13th Afternoon - Jamie with supervision of Suresh)

Cabling2: Done - RF generation box (IOO Rack) to the demodulator

Demodulator: In Progress - Test and install (13th night - Kiwamu with supervision of Suresh)

LSC model: Done - Run the new LSC model. (It is named as "C1LST" so far) (13th evening - Jamie)

LSC medm: Done on 14th - Modify the current LSC medm screens Update the EPICS database Adjust the matrices (- Jenne with supervision of Koji)

  4522   Thu Apr 14 00:21:28 2011 KojiUpdateCDSNew C1LSC code running

[Jamie, Jenne, Koji]

We installed the new c1lsc and started the process.

We still need to configure bunch of the EPICS variables, matrices, and some of the filters.
This should be done in order to transmit the signals to the suspensions.
Jenne is going to work on this task tomorrow (Friday) morning,
and Koji will take over the task afternoon/evening.

  4521   Wed Apr 13 23:32:07 2011 Aidan, JamieConfigurationLSCAS PD and Camera installed

I spent some time tracking down the AS beam which had vanished from the AP table. Eventually, by dramatically mis-aligning SRM, PRM and ITMY, returning BS to its Jan 1st PITCH and YAW values and tweaking the ITMX alignment [actual values to follow], I was able to get an AS beam out onto the AP table. I verified that it was the prompt reflection off ITMX by watching it move as I changed the YAW of that optic and watching it stay stationary as I changed the YAW of ITMY.

Jamie and I then steered the beam through a 2" PLCX-50.8-360.6 lens and placed the RF PD (AS55) at the focus. Additionally, we installed the AS camera to observe the leakage field through a Y1S steering mirror (as shown in the attached diagram).

Currently the PD has power but the RF and DC outputs are not connected to anything at the moment.

Atm 2 by Steve

 

 

Attachment 1: AS_beam.jpg
AS_beam.jpg
Attachment 2: P1070546.JPG
P1070546.JPG
  4520   Wed Apr 13 16:56:08 2011 BryanConfigurationGreen LockingY-ARM Green-Locked!

 Locked!

The Y-arm can now be locked with green light using the universal PDH servo. Modulation frequency is now 277kHz - chosen because it seems to produce smaller offsets due to AM effects

To lock, turn on the servo, align the system to give nice circular-looking TEM_00 resonances, and wait for a good one. It'll lock on a decent mode for a few seconds and then you can turn on the local boost and watch it lock for minutes and minutes and minutes.

The suspensions are bouncing around a bit on the Y-arm and the spot is quite low on the ETMY and a little low on ITMY, but from this point it can be tweaked and optimised.

 

 

 

  4519   Wed Apr 13 16:38:17 2011 Larisa ThorneUpdateElectronicsVideo MUX camera/monitor check

 [Kiwamu, Larisa]

 

The following Video MUX inputs(cameras) and outputs(monitors) have been checked:

MC2F, FI, AS Spare, ITMYF, ITMXF, ETMYF, ETMXF, PSL Spare, ETMXT, MC2T, POP, MC1F/MC3F, SRMF, ETMYT, PRM/BS, CRT1(MON1), ETMY Monitor, CRT2(MON2), CRT4(MON4), MC1 Monitor, CRT3(MON3), PSL1 Monitor, PSL2 Monitor, CRT6(MON6), CRT5(MON5), ETMX Monitor, MC2 Monitor, CRT9, CRT7(MON7), CRT10, and Projector.

 

Their respective statuses have been updated on the wiki:   (wiki is down at the moment, I will come back and add the link when it's back up)

  4518   Wed Apr 13 11:34:07 2011 josephbUpdateCDSFixed IFO_ALIGN.adl

Problem:

I switched the ITMX and ITMY control channels yesterday, but forgot to update the IFO_ALIGN.adl file (/opt/rtcds/caltech/c1/medm/c1ifo/) which had the control labels swapped to make life easier.

Solution:

I swapped ITMX and ITMY control locations on the screen.

Question:

Are there any other screens involving ITMX and ITMY that had controls reversed to make life easier?

  4517   Tue Apr 12 18:15:07 2011 kiwamuSummaryIOORF combiner is more like attenuator

I realized that my impedance matching theory on an RF combiner was wrong !

In fact an RF combiner behaves more like an attenuator according to a reflection measurement that I did today.

A 3-way combiner reduces power of an input signal by a factor of 4.8 dB because it can be also considered as a 3-way splitter.

So it is just a lossy component or in other words it is just an attenuator.

 

(reflection measurement)

To check my speculation that I posted on #4504 I measured reflection coefficients for both cases.

In the measurement I used a heliax cable, which goes from 1X2 rack to the PSL table with a length of about 10 m. Note that this is the cable that had been used as '33 MHz EOM'.

At the input of the heliax cable it was connected to a direction coupler to pick off reflections and the reflected signal was sampled in AG4395A.

The other end of the cable (output side of the cable) was basically connected to the resonant box.

Then I did a reflection measurement for both cases as drawn in this entry (see #4504).

  - case 1 -  the combiner was inserted at the input side of the heliax cable.

  - case 2 - the combiner was directly attached to the resonant box

On the combiner, ZFSC-3-13, the port 1 and 2 were terminated with 50 Ohm, therefore the port 3 was used as an input and the source port is the output.

Here is a resultant plot of the reflection measurements.

reflection_test.png

Note that whole data are calibrated so that it gives 0 dB when the output side of the heliax is open.

There are two things we can notice from this plot:

 (1) The reflection coefficient at the resonant frequencies (where notches appear) are the same for both cases.

 (2) Over the measured frequency range the reflections were attenuated by a factor of about 9.6 dB , which is twice as large as the insertion loss of the combiner.

These facts basically indicates that  the RF combiner behaves as a 4.8 dB attenuator.

Hence the location of the combiner doesn't change the situation in terms of RF reflections.

Quote from #4505

 The input impedance of the resonant box was measured when an RF combiner was attached to the box.

Indeed the combiner makes the impedance more 50 Ohm and reduces the reflection.

 

 

  4516   Tue Apr 12 16:01:33 2011 josephbUpdateGeneralRFM errors

Problem:

Currently the c1scy, c1mcs, and c1rfm models are reporting an error with receiving some data sent over the GE Fanuc Reflected memory cards.

To be more exact, the C1:SUS-ETMY_ALS signal from the c1gcv FE code on the c1ioo computer going too the Y end is not being received However, the C1:SUS-ETMY_LSC signal is.  So the physical RFM card seems to be working.

Similarly, the TRY signal is being sent correctly from the Y end computer.  The X end is working fine and receiving both LSC and ALS signals.

The c1mcs and c1rfm models also receive data from the c1ioo computer and reporting receiving errors.

Theory:

Because the RFM cards are transmitting and receiving at least some channels, I'm guessing there was changes made to the C1.ipc file, which defines the memory locations of these various channels on the RFM network, and that when a model was rebuilt, a different one using the previous IPC file was not, and thus one of the computer is going to the wrong place to either read or write data.

Tomorrow, I'm planning on the  following:

1) Clean out the C1.ipc file (/opt/rtcds/caltech/c1/chans/ipc/)

2) Rebuild all models

3) Run activate_daq.py script

4) Restart models via script

If this doesn't clear up the problem, I'll continue  to bug hunt.

  4515   Tue Apr 12 12:01:30 2011 josephbUpdateGeneralIFO controls, now with 10% less lying (ITMX/ITMY controls swapped)

The ITMX/ITMY control swap is complete.

The steps from this elog were followed.

In addition, I did a burt restore of c1sus, c1mcs.

I then swapped all the gain settings from ITMX to ITMY, and reenabled the watchdogs.

I did some basic kick tests (1000 counts into UL coil) and confirmed channels like C1:SUS-ITMX_ULPD_VAR (watchdogs mV readback) corresponded to the correct optic.  I also checked that the POS, PIT, YAW, SIDE produced reasonable damping when engaged.

ELOG V3.1.3-